OpenStudy (gorica):

A=[1, -5, 9; 2, 0, -1; 1, 2 , -4] A^20=? using Cayley–Hamilton theorem!

OpenStudy (anonymous):

did you compute the characteristic polynomial? maybe we can use that, and some gimmick to find \(A^{20}\)

OpenStudy (anonymous):

looks like it is \[P(x)=-x^3-3 x^2+x+3\] then \(-A^3-3A^2+A+3=0\) and so \[A^3=-3A^2+A+3\] then multiply both sides by \(A\) and get \[A^4=-3A^3+A^2+3A+3\] now replace \(A^3\) in this expression by \(-3A^2+A+3\) and get \[A^4=-3(-3A^2+A+3)+A^2+3A+3=10A^2-6\]

OpenStudy (anonymous):

well that took a while, but maybe now there is some other trick to find \(A^{20}\) without continuing along this line of work, but i am pretty sure this is one way to do it

OpenStudy (anonymous):

the point being that by repeating this process you will be able to compute \(A^{20}\) as some quadratic polynomial in \(A\)

OpenStudy (anonymous):

now maybe multiply both sides by \(A^2\) and get \[A^6=10A^4-6A^2=10(10A^2-6)-6A^2=96A^2-60\] hmm this is going to be ugly, maybe my algebra is wrong

OpenStudy (gorica):

I have this problem solved. From characteristic polynomial \[A ^{3}=-3A ^{2}+A+3E\] and then here is written \[A ^{n}=a _{n}A ^{2}+bnA+c _{n}E\] and after computing by A I have \[A ^{n+1}=a _{n}A ^{3}+b _{n}A ^{2}+c _{n}A\]\[A ^{n+1}=(b _{n}-3a _{n})A ^{2}+(b _{n}+c _{n})A+3a _{n}E\] \[a _{n+1}=b _{n}-3a _{n} \rightarrow a _{n+1}=a _{n-1}+3a _{n-2}-3a _{n} \] \[b _{n+1}=a _{n}+c _{n} \rightarrow b _{n+1}=a _{n}+3a_{n-1}\] \[c _{n+1}=3a _{n}\] now, using \[A ^{n}=a _{n}A ^{2}+b _{n}A+c _{n}E\] I have to compute coefficients for n=0, 1, ..., 20 I don't know how did they get those coefficients above

OpenStudy (gorica):

*multiplying by A

OpenStudy (anonymous):

which coefficients?

OpenStudy (anonymous):

we have them for \(n=3\) because \(A^3=-3A^2+A+3E\) so \[a_3=-3,b_3=1,c_3=3\]

OpenStudy (gorica):

an+1=bn−3an→an+1=an−1+3an−2−3an bn+1=an+cn→bn+1=an+3an−1 cn+1=3an these recurrent formulas, I don't know how did they get them

OpenStudy (anonymous):

oh ok lets go step by step is this line \[A ^{3}=-3A ^{2}+A+3E\] clear?

OpenStudy (gorica):

yes :D

OpenStudy (anonymous):

did you see above how i write \(A^4\) as a quadratic in \(A\)?

OpenStudy (gorica):

I did

OpenStudy (anonymous):

is it clear then that \(A^n\) can be written as a quadratic in \(A\)

OpenStudy (gorica):

yes

OpenStudy (anonymous):

then this line is clear right? \[A ^{n+1}=a _{n}A ^{3}+b _{n}A ^{2}+c _{n}A\]

OpenStudy (anonymous):

oh i meant \[A^n=a_nA^2+b_nA+c_nE\] that says \(A^n\) is a quadratic in \(A\)

OpenStudy (anonymous):

multiply by \(A\) and get \[A ^{n+1}=a _{n}A ^{3}+b _{n}A ^{2}+c _{n}A\]

OpenStudy (anonymous):

that part is ok or no?

OpenStudy (gorica):

yes, that's ok

OpenStudy (anonymous):

ok now we have \[A^{n+1}=a_{n+1}A^2+b_{n+1}A+c_{n+1}E=a_nA^3+b_nA^2+c_nA\]

OpenStudy (anonymous):

i.e. two expressions for \(A^{n+1}\) is that ok?

OpenStudy (gorica):

yes

OpenStudy (anonymous):

so now to relate the coefficients, you can replace \(A^3\) on the right hand side by \[-3A^2+A+3E\] analagous to how i wrote \(A^4\) in term of \(A^2\)

OpenStudy (anonymous):

i am going to make a guess that it is this step that was not obvious, since the just wrote the result of that computation, rather than the actual computation i could be wrong

OpenStudy (gorica):

Thank you, I got it :) Thank you very much :)

OpenStudy (anonymous):

yw glad to help